OpenStudy (anonymous):

negative 10 equals x minus 21

OpenStudy (studygurl14):

-10 = x -21 Add 21 to both sides 11 = x

OpenStudy (anonymous):

wouldnt it be negative 11?

OpenStudy (studygurl14):

No it wouldn't because -10 + 21 = 11

OpenStudy (anonymous):

ohh, no its positive i see it, thanks

OpenStudy (studygurl14):

you're welcome :)